OpenStudy (anonymous):

I need your help, PLEASE! What is the solution for 4x − 4 > 12?

OpenStudy (anonymous):

4X-4=12 4X=12+4 4X=16 X=16/4 X=4

OpenStudy (anonymous):

4x-4>12 4x>12+4 4x>16 x>4

OpenStudy (anonymous):

damn i forgot the >

OpenStudy (anonymous):

thank you guys :)

OpenStudy (anonymous):

solution is (4,infinity)